# A Simple First-Order Algorithm for Full-Rank Equality Constrained Optimization

### v1

@@ Abstract
A very simple first-order algorithm is proposed for solving nonlinear optimization problems with nonlinear equality constraints. This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility, without using a merit function or filter. The tangent steps are based on the AdaGrad method for unconstrained minimization. The objective function is never evaluated by the algorithm, making it suitable for noisy problems. Its worst-case evaluation complexity is analyzed, yielding a global convergence rate in ${\pazocal{O}}(1/\sqrt{k})$, which matches the best known rate of first-order methods for unconstrained problems. Numerical experiments are presented suggesting that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems, and that its reliability is remarkably stable in the presence of noise on the gradient.
Keywords: Equality constrained optimization, objective-function-free optimization (OFFO), first-order methods, AdaGrad, evaluation complexity.
@@ 1 Introduction
Solving nonlinear problems with nonlinear equality constraints is a central part of more general constrained optimization, and the purpose of this paper is to contribute to the subject by proposing a very simple algorithm for the case where the constraints’ Jacobian is full-rank. We introduce a method drawing its inspiration both from the “trust-funnel” approaches [17, 8] and the objective-function free (OFFO) first-order algorithms that have become very popular because of their successful use in deep learning applications. Trust-funnel methods themselves exploit the older idea of decomposing the step computed within an optimization algorithm for constrained problems into a “tangential” and a “normal” components. Advocated in particular by Byrd and Omojokun [24] for sequential quadratic programming, this last idea is to consider that a useful minimization step should, on the one hand, improve the value of the objective function without deteriorating the constraints’s violation too much (this is the tangential step, because it can be interpreted as a step (nearly) tangent to the constraint’s surface) and, on the other hand, improve feasibility (the normal step, mostly orthogonal to the constraints’ surface). This technique has been successfully used by many algorithms, and in particular by trust-funnel ones. In this case, the two types of steps are treated separately and a choice between normal and tangential steps is made at every iteration by comparing constraint violation (primal feasibility) with some dual optimality measure. Our new proposed algorithm uses a similar mechanism, although considerably simplified. Our second source of inspiration is the thriving field of first-order methods for unconstrained problems, and, more specifically, the well-known and well understood AdaGrad [10, 29, 9, 19] algorithm. Such methods are attractive because they avoid computing the value of the objective function entirely (they only rely on gradients). This feature makes them quite robust in the presence of noise on the gradient, as is commonly the case when its evaluation involves subsampling. Again, our new method exploits this OFFO [^1] Objective-Function-Free Optimization. strategy to achieve robustness. In effect, it consists of a simplified trust-funnel method in which the tangential step is computed using the AdaGrad algorithm.
@@
Other OFFO methods for constrained optimization have been proposed. The stochastic algorithm proposed in [1] uses a combined tangential and normal step for a quadratic model of a merit function. A (nontrivial) stochastic analysis of its convergence (albeit not its worst-case complexity) is provided. It also allows the use of second-order information, at the cost of a somewhat complicated algorithm. [7] extends this proposal to handle inequality constraints and provides an updated convergence analysis. The algorithm of [11] is more similar to ours, but differs in significant aspects. Although it uses tangential and normal steps, it does so at every iteration. The step in the Jacobian’s nullspace is computed by a standard trust-region technique and a special rule is provided to update its radius. It also relies on an explicit penalty parameter with is updated in the course of the algorithm. Convergence of the algorithm is also analyzed in a stochastic setting. The method of [4] is inspired by filter methods (another class of methods for nonlinear optimization [13, 14, 12]) and also alternates between steps which attempts to reduce constraints’ infeasibilities and “optimality” steps improving the objective function value, albeit ignoring the presence of constraints in the second case. Good numerical results are reported on deep learning problems, but no convergence theory or complexity analysis is provided [^2] Since the optimality steps ignore the constraints’ geometry, it seems possible that their effect might significantly decrease feasibility and that cycling could occur..
@@
    We provide a complete worst-case complexity analysis for the case where the objective’s gradient and the constraints are deterministic, under assumptions similar to those used in [1, 7].
@@
    We illustrate the practical behaviour of the method on a subset of the CUTEst problems [15] as provided by the S2MPJ environment [21].
@@
2 An adaptative switching algorithm
@@
using the purely first-order ADSWITCH method, which we now describe after establishing some notation. In (2.1), $f$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into IR and $c$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into $\hbox{I\thinspace R}^{m}$ ($m\leq n$). For a given vector $x$, we assume that we can compute the gradient $g(x)=\nabla_{x}f(x)$, as well as the (exact) Jacobian of the constraints at $x$
@@
Assuming $J(x)$ has full rank, we may also compute the orthogonal projection onto its nullspace by
@@
Given this projection, the projected gradient is then
$g_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}P_{T}(x)\,g(x)$ (2.2)
@@
and, for $\rho$ fixed, the associated “augmented-Lagrangian-like” Lyapunov function
@@
$g_{T}(x)=\nabla f(x)+J(x)^{T}\widehat{\lambda}(x)\;\;\mbox{ and }\;\;J(x)\,g_{T}(x)=0,$ (2.7)
@@
Step 0: Initialization. A starting point $x_{0}$ is given, together with constants $\omega>0$, $\theta>1$, $\beta,\eta,\varsigma\in(0,1]$. Set $\Gamma_{0}=0$ and $k=0$. Step 1: Evaluations. Evaluate $c_{k}=c(x_{k})$, $J_{k}=J(x_{k})$ and $g_{k}=\nabla_{x}f(x_{k})$.
@@
Step 2: Tangential step: If $\|c_{k}\|\leq\beta\alpha_{T,k}\|g_{T,k}\|$ (2.12) then set $x_{k+1}=x_{k}-\alpha_{T,k}g_{T,k}$ (2.13) and $\Gamma_{k+1}=\Gamma_{k}^{+},$ (2.14) and go to Step 4. Step 3: Normal step: Otherwise, find a step $s_{N,k}$ in the range space of $J_{k}^{T}$ such that $\|s_{N,k}\|\leq\theta\|c_{k}\|$ (2.15) and there exists $\kappa_{\mbox{\tiny nrm}}>0$ independent of $k$ such that ${\scriptstyle\frac{1}{2}}\|c(x_{k}+s_{N,k})\|^{2}\leq{\scriptstyle\frac{1}{2}}\|c_{k}\|^{2}-\kappa_{\mbox{\tiny nrm}}\|J_{k}^{T}c_{k}\|^{2},$ (2.16) and set $x_{k+1}=x_{k}+s_{N,k}.$ (2.17) Step 4: Loop. Increment $k$ by one and go to Step 2.
@@
  • The statement of the procedure to find the normal step in Step 3 may seem abstract because $\kappa_{\mbox{\tiny nrm}}$ may not be known, but, fortunately, there exist several standard techniques to achieve (2.15)-(2.16).
    A first technique is to use a standard Armijo backtracking linesearch along the steepest-descent direction $J_{k}^{T}c_{k}$ (see Lemma 3.3). It is well-known that (2.16) holds in this case with $\kappa_{\mbox{\tiny nrm}}=1/L_{JTc}$ (see [3, Lemma 2.2.1], for instance). More generally, any “gradient related” direction $d_{k}$, that is any $d_{k}$ satisfying (2.15) and $d_{k}^{T}J_{k}^{T}c_{k}\leq-\kappa_{\mbox{\tiny grl}}\|J_{k}^{T}c_{k}\|^{2}$ for some $\kappa_{\mbox{\tiny grl}}>0$, can be used to initiate a backtracking linesearch.
    A second possibility follows along this line and is to select a positive definite matrix $B_{k}$ with bounded condition number, to define $d_{k}=-B_{k}^{-1}J_{k}^{T}c_{k}$ and then to perform a backtracking linesearch along $d_{k}$ (see [5, Section 10.3.1] for a proof that $d_{k}$ is a gradient-related direction), in which case a similar result holds. For instance, using $B_{k}=J_{k}^{T}J_{k}+\delta I_{n}$ for some $\delta\geq 0$ yields the regularized Gauss-Newton step
@@
    Suitable variants of trust-region steps are also possible [5, Section 10.3.2], themselves involving several subvariants such as the truncated conjugate gradient [26, 27], CGLS [22] or LSQR [25]. Because iterates generated by these methods are linear combinations of the gradients of ${\scriptstyle\frac{1}{2}}\|c_{k}+J_{k}s\|^{2}$, this guarantees that the final step belongs to the range of $J_{k}^{T}$.
@@
    $s_{N,k}=\left\{\begin{array}[]{ll}-Q_{k,[1:n,1:r]}R_{k,[1:r,1:r]}^{-T}c_{k}&\;\;\mbox{if }\;\;\delta=0,\\ -Q_{k,[1:n,1:m]}R_{k,[1:m,1:m]}(R_{k,[1:m,1:m]}^{T}R_{k,[1:m,1:m]}+\delta I_{m})^{-1}c_{k}&\;\;\mbox{otherwise.}\end{array}\right$. (2.19)
  • Observe that the ADSWITCH algorithm does not involve any explicit merit function or filter to control overall progress, but rely on the simple switching condition (2.12) (hence its name). As we will see below, an implicit merit function (a modified augmented-Lagrangian with suitably chosen parameters) is however crucial for our theoretical argument.
  • ADSWITCH differs in several aspects from the trust-funnel methods of [17, 8]. The first is that it avoids evaluating the objective function in order to improve its reliability in the presence of noise. The second is that it also avoids estimating Lagrange multipliers, but at the cost of computing the exact projection $P_{T}(\cdot)$. The third is that the trust-region approach, which is explicit in trust-funnel methods, is only implicit in ADSWITCH. Indeed it has been argued in [18] that the (unconstrained) AdaGrad algorithm can be viewed as a trust-region method with a specific radius management. Moreover, (2.15) can also be viewed as a trust-region constraint (such a condition is also imposed in [17]), but without the need to explicitly update its radius. Finally, ADSWITCH [^3] In its present incarnation. is more limited than trust-funnel methods in that it does not handle approximate projections in the tangent plane.
In what follows, we denote by $\{k_{\tau}\}\subseteq\{k\}$ the index subsequence of “tangential iterations”, that is iterations where (2.12) holds. We also denote $\{k_{\nu}\}=\{k\}\setminus\{k_{\tau}\}$ the index subsequence of “normal iterations”.
@@ 3.1 Assumptions
For all $x\in\hbox{I\thinspace R}^{n}$, $\|g(x)\|\leq\kappa_{g}$ where $\kappa_{g}\geq\eta\beta$.
@@
For all $x\in\hbox{I\thinspace R}^{n}$, $\sigma_{\min}(J(x))\geq\sigma_{0}\in(0,1]$.
@@
Note that these assumptions are very similar to those of [1, 7]. In particular, AS.5 is also part of Assumption 1 in this last reference.
We finally assume, without loss of generality [^4] The parameter $\varsigma$ may be re-adjusted downwards at the first tangential iteration, if necessary., that

### v2

@@ Abstract
A very simple first-order algorithm is proposed for solving potentially stochastic nonlinear optimization problems with deterministic nonlinear equality constraints. This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility, without using a merit function or filter. The tangent steps are based on the AdaGrad method for unconstrained minimization. The objective function is never evaluated by the algorithm, making it suitable for noisy problems. Its worst-case evaluation complexity and convergence is analyzed both in the deterministic and stochastic cases, yielding a global convergence rate in ${\pazocal{O}}(1/\sqrt{k})$ for the deterministic case and in ${\pazocal{O}}(1/k^{1/4})$ for the stochastic one, which match the best known rates of first order methods for unconstrained problems. Numerical experiments are presented suggesting that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems, and that its reliability is remarkably stable in the presence of noise on the gradient.
Keywords: Equality constrained optimization, objective-function-free optimization (OFFO), first-order methods, AdaGrad, evaluation complexity, stochastic analysis.
@@ 1 Introduction
Solving nonlinear problems with nonlinear equality constraints is a central part of more general constrained optimization, and the purpose of this paper is to contribute to the subject by proposing a very simple algorithm for the case where the constraints’ Jacobian is full-rank. We introduce a method drawing its inspiration both from the “trust-funnel” approaches [15, 7] and the objective-function free (OFFO) first-order algorithms that have become very popular because of the successful use in deep learning applications. Trust-funnel methods themselves exploit the older idea of decomposing the step computed within an optimization algorithm for constrained problems into a “tangential” and a “normal” components. Advocated in particular by Byrd and Omojokun [22] for sequential quadratic programming, the last idea is to consider that a useful minimization step should, one one hand, improve the value of the objective function without deteriorating the constraints’s violation too much (this is the tangential step, because it can be interpreted as a step (nearly) tangent to the constraint’s surface) and, on the other hand, improve feasibility (the normal step, mostly orthogonal to the constraints’ surface). This technique has been successfully used by many algorithms, and in particular by trust-funnel ones. In this case, the two types of steps are treated separately and a choice between normal and tangential steps is made at every iteration by comparing constraint violation (primal feasibility) with some dual optimality measure. Our new proposed algorithm uses a similar mechanism, although considerably simplified. Our second source of inspiration is the thriving field of first-order methods for unconstrained problems, and, more specifically, the well-known and well understood AdaGrad [9, 27, 8, 17] algorithm. Such methods are attractive because they avoid computing the value of the objective function entirely (they only rely on gradients). This feature makes them quite robust in the presence of noise on the gradient, as is commonly the case its evluation involves subsampling. Again, our new method exploits this OFFO [^1] Objectibe-Function-Free Optimization. strategy to achieve robustness. In effect, it consists of a simplified trust-funnel method in which the tangential step is computed using the AdaGrad algorithm.
@@
Other OFFO methods for constrained optimization have been proposed. The stochastic SQP method of [6] is more general in that it also handles inequality constraints and provides a (nontrivial) stochastic analysis of its convergence (albeit not its worst-case complexity). It also allow the use of second-order information, at the cost of a quite complicated algorithm. The method of [3] is inspired by filter methods (another class of methods for nonlinear optimization [11, 12, 10]) and also alternates between steps which attempts to reduce constraints’s infeasibilities and “optimality” steps improving the objective function value, albeit ignoring the presence of constraints in the second case. Good numerical results are reported on deep learning problems, but no convergence theory or complexity analysis is provided [^2] Since the optimality steps ignore the constraints’s geometry, it seems possible that their effect might significantly decrease feasibility and that cycling could occur..
@@
    We provide a complete worst-case complexity analysis for the case where the objective’s gradient can be deterministic or contaminated by random noise and the constraints are deterministic, under assumptions similar to those used in [6].
@@
    We illustrate the practical behaviour of the method on a subset of the CUTEst problems [13] as provided by the S2MPJ environment [19], both in the deterministic and noisy contexts.
@@
2 An adapative switching algorithm
@@
using the purely first-order ADSWITCH method, which we now describe after establishing some notation. In (2.1), $f$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into IR and $c$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into $\hbox{I\thinspace R}^{m}$ ($m\leq n$). For a given vector $x$, we assume that we can compute a random approximation $g(x)$ of the gradient $G(x)=\nabla_{x}f(x)$, as well ase the (exact) Jacobian of the constraints at $x$
@@
and the orthogonal projection onto its nullspace by
@@
Given this projection, the projected gradient and its approximation are then
$G_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}P_{T}(x)\,G(x)\;\;\mbox{ and }\;\;g_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}P_{T}(x)\,g(x)$ (2.2)
@@
and, for $\rho$ fixed, the associated “augmented-Lagrangian-like” Lyapounov function
@@
$G_{T}(x)=\nabla f(x)+J(x)^{T}\widehat{\lambda}(x)\;\;\mbox{ and }\;\;J(x)\,g_{T}(x)=J(x)\,G_{T}(x)=0,$ (2.7)
@@
Step 0: Initialization. A starting point $x_{0}$ is given, together with constants $\omega>0$, $\theta>1$, $\beta,\eta,\varsigma\in(0,1]$. Set $\Gamma_{-1}=0$ and $k=0$. Step 1: Evaluations. Evaluate $c_{k}=c(x_{k})$, $J_{k}=J(x_{k})$ and $g_{k}\approx\nabla_{x}f(x_{k})$.
@@
Step2: Tangential step: If $\|c_{k}\|\leq\beta\alpha_{T,k}\|g_{T,k}\|$ (2.12) then set $x_{k+1}=x_{k}-\alpha_{T,k}g_{T,k}$ (2.13) and $\Gamma_{k+1}=\Gamma_{k}^{+},$ (2.14) and go to Step 4. Step 3: Normal step: Otherwise, find a step $s_{N,k}$ in the range space of $J_{k}^{T}$ such that $\|s_{N,k}\|\leq\theta\|J_{k}^{T}c_{k}\|$ (2.15) and there exists $\kappa_{\mbox{\tiny nrm}}>0$ independent of $k$ such that ${\scriptstyle\frac{1}{2}}\|c(x_{k}+s_{N,k})\|^{2}\leq{\scriptstyle\frac{1}{2}}\|c_{k}\|^{2}-\kappa_{\mbox{\tiny nrm}}\|J_{k}^{T}c_{k}\|^{2},$ (2.16) and set $x_{k+1}=x_{k}+s_{N,k}.$ (2.17) Step 4: Loop. Increment $k$ by one and go to Step 2.
@@
  • The statement of the procedure to find the normal step in Step 3 may seem abstract because $\kappa_{\mbox{\tiny nrm}}$ may not be known, but, fortunately, there exist several standard techniques to achieve (2.16).
    A first technique is to use a standard Armijo backtracking linesearch along the the steepest-descent direction $J_{k}^{T}c_{k}$ with initial step size $\theta$. It is well-known that (2.16) holds in this case with $\kappa_{\mbox{\tiny nrm}}=1/L_{JTc}$ (see [2, Lemma 2.2.1], for instance). More generally, any “gradient related” direction $d_{k}$, that is any $d_{k}$ satisfying (2.15) and $d_{k}^{T}J_{k}^{T}c_{k}\leq-\kappa_{\mbox{\tiny grl}}\|J_{k}^{T}c_{k}\|^{2}$ for some $\kappa_{\mbox{\tiny grl}}>0$, can be used to initiate a backtracking linesearch.
    A second possibility follows along this line and is to select a positive definite matrix $B_{k}$ with bounded condition number, to define $d_{k}=-B_{k}^{-1}J_{k}^{T}c_{k}$ and then to perform a backtracking linesearch along $d_{k}$ (see [4, Section 10.3.1] for a proof that $d_{k}$ is a gradient-related direction), in which case a similar result holds. For instance, using $B_{k}=J_{k}^{T}J_{k}+\delta I_{n}$ for some $\delta\geq 0$ yields the regularized Gauss-Newton step
@@
    Suitable variants of trust-region steps are also possible [4, Section 10.3.2], themselves involving several subvariants such as the truncated conjugate gradient [24, 25], CGLS [20] or LSQR [23]. Because iterates generated by these methods are linear combinations of the gradients of ${\scriptstyle\frac{1}{2}}\|c_{k}+J_{k}s\|^{2}$, this guarantees that the final step belongs to the range of $J_{k}^{T}$.
@@
    $s_{N,k}=\left\{\begin{array}[]{ll}-Q_{k[1:n,1:r]}R_{k[1:r,1:r]}^{-T}c_{k}&\;\;\mbox{if }\;\;\delta=0,\\ -Q_{k[1:n,1:m}R_{k[1:m,1:m]}(R_{k[1:m,1:m]}^{T}R_{k[1:m,1:m]}+\delta I_{m})^{-1}c_{k}&\;\;\mbox{otherwise.}\end{array}\right$. (2.19)
  • Observe that the ADSWITCH algorithm does not involve any explicit merit function or filter to control overal progress, but rely on the simple switching condition (2.12) (hence is name). As we will see below, an implicit merit function (an modified augmented-Lagrangian with suitably chosen parameters) is however crucial for our theoretical argument.
  • ADSWITCH differs in several aspects from to the trust-funnel methods of [15, 7]. The first is that it avoids evaluating the objective function in order to improve its reliability in the presence of noise. The second is that it also avoids estimating Lagrange multipliers, but at the cost of computing the exact projection $P_{T}(\cdot)$. The third is that the trust-region approach, which is explicit in trust-funnel methods, is only implicit in ADSWITCH. Indeed it has been argued in [16] that the (unconstrained) AdaGrad algorithm can be viewed as a trust-region method with a specific radius management. Moreover, (2.15) can also be viewed as a trust-region constraint, but without the need to explicitly update its radius. Finally, ADSWITCH [^3] In its present incarnation. is more limited that trust-funnel methods in that it does not handle approximate projections in the tangent plane.
In what follows, we denote by $\{k_{\tau}\}\subseteq\{k\}$ the index subsequence of tangential iterations, that is iterations where (2.12) holds. We alse denote $\{k_{\nu}\}=\{k\}\setminus\{k_{\tau}\}$ the index subsequence of normal iterations. We will use the abbreviations
$\mathbb{E}^{\tau}_{k}\!\left[\cdot\right]\stackrel{{\scriptstyle\rm def}}{{=}}\mathbb{E}_{k}\!\left[\cdot|k\in\{k_{\tau}\}\right]\;\;\mbox{ and }\;\;\mathbb{E}^{\nu}_{k}\!\left[\cdot\right]\stackrel{{\scriptstyle\rm def}}{{=}}\mathbb{E}_{k}\!\left[\cdot|k\in\{k_{\nu}\}\right].$
@@ 3.1 Assumptions
For all $k\geq 0$, $\|G_{k}\|\leq\kappa_{g}$ where $\kappa_{g}\geq\eta^{2}\beta^{2}$.
@@
For all $k\geq 0$, $\sigma_{\min}(J_{k})\geq\sigma_{0}\in(0,1]$.
@@
Note that these assumptions are very similar to those of [6]. In particular, AS.5 is also part of Assumption 1 in this reference.
Because of the random nature of the gradient estimator, the ADSWITCH algorithm generates a random process where, for a given iterate $x_{k}$, the oracle computes the gradient oracle $g_{k}=g(x_{k},\xi)$ where $\xi$ is a random variable (whose distribution may depend on $x_{k}$), with probability space $(\Omega,{\pazocal{F}},\mathbb{P})$. The expectation conditioned to knowing $g_{0},\ldots,g_{k-1}$ will be denoted by the symbol $\mathbb{E}_{k}\!\left[\cdot\right]$. Note that the projection $P_{T}(x_{k})$ is measurable with respect to the past, and so is $\widehat{\lambda}(x_{k})$. We complete our assumption as follows.
AS.8 :
For all $k\geq 0$, the estimator of $g_{k}$ is unbiased, that is $\mathbb{E}_{k}\!\left[g_{k}\right]=G_{k}$.
AS.9 :
For all $k\geq 0$, $\|g_{k}\|\leq\kappa_{g}$ where $\kappa_{g}\geq\eta^{2}\beta^{2}$.
Observe that, because $P_{T}(x_{k})$ is a linear operator,
$\mathbb{E}_{k}\!\left[g_{T,k}\right]=\mathbb{E}_{k}\!\left[P_{T}(x_{k})g_{k}\right]=P_{T}(x_{k})\mathbb{E}_{k}\!\left[g_{k}\right]=P_{T}(x_{k})G_{k}=G_{T,k}.$ (3.1)
Using Jensen’s inequality and the convexity of the norm, we also have that
$\mathbb{E}\!\left[\|G_{T,k}\|\right]=\mathbb{E}\!\left[\|\mathbb{E}_{k}\!\left[g_{T,k}\right]\|\right]\leq\mathbb{E}\!\left[\mathbb{E}_{k}\!\left[\|g_{T,k}\|\right]\right]=\mathbb{E}\!\left[\|g_{T,k}\|\right].$ (3.2)
We next require a “directional root mean square” condition [1] along the tangential step $\alpha_{T,k_{\tau}}g_{T,k_{\tau}}$ given by
AS.10:
There exists a contant $\kappa_{\mbox{\tiny dir}}>0$ such that, for all $k\geq 0$,
$\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}|(G_{T,k}-g_{T,k})^{T}g_{T,k}|\right]\leq\frac{\kappa_{\mbox{\tiny dir}}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
We finally assume, without loss of generality [^4] The parameter $\varsigma$ may be re-adjusted downwards at the first tangential iteration, if necessary., that, for all realizations,